Calculus I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 15

Assigned Dec 22, 2011.

1. Section 8.2: Problems: 7, 9, 13, 19, 25, 27, 29, 30, 43, 47. Hint for Problem 29: read Example 6.

2. Section 8.3: Problems: 11, 15, 17, 19, 33, 35, 39, 41, 42, 43. Hint for Problem 41, 42: break the square root and cubic root by a proper change of variable.